An ideal gas is expanding such that PT = constant. The coefficient of volume expansion of the gas is
 

Grade:12th pass

2 Answers

Arun
6 years ago
 
Dear student
 
PT = k given AND the coeff. of volume expansion of the gas is given as:
Y =(1/V)(dV/dT)
From ideal gas eqn.,
PV = nRT
or
kV = nRT^2
dV/dT => (2T)*(nR/k)
and
Y = 2nRT/kV
or
2PV/kV => 2P/k => 2P/PT => 2/T
 
 
Regards
Arun (askIITians forum expert)
Yash Chourasiya
3 years ago
Dear Student

Given PT = k

coeff. of volume expansion of the gas is given as (γ) =(1/V)(dV/dT)
PV = nRT
or
kV = nRT2
dV/dT => (2T)*(nR/k)
and
γ= 2nRT/kV
or
2PV/kV => 2P/k => 2P/PT
=> 2/T

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya